OpenStudy (calculusxy):

Piper works at a camera store. He is paid an hourly rate plus 16% commission on everything he sells One week, he was paid $365 for working 20 hours and selling $1,500 worth of camera equipment. What is his hourly rate?

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

earned = 20 hours * rate + commission on 1500

OpenStudy (calculusxy):

So would the earned include $365 + $1,500 ?

OpenStudy (phi):

no. commission means if you sells cameras, and at the end of the week, the customers paid 1500 for all the cameras they bought from him, he gets a small share (he does not get it all, because the store has to buy cameras from the factory, and pay for renting the store, and making a profit.)

OpenStudy (calculusxy):

so that means that will the earned be equal to 365 + 1500(0.16)?

OpenStudy (phi):

the 365 is how much he got from both the commission and for rate * 20 hours

OpenStudy (phi):

He makes most of his money from the commission, and a small amount from the hourly pay: hours * rate + commission = 365

OpenStudy (phi):

You know hours is 20 you know the commission is 1500*0.16 you need to figure out rate

OpenStudy (calculusxy):

So is this something to do with isolating a variable to find out the rate?

OpenStudy (phi):

yes. First write the equation 20x + 1500*0.16 = 365 solve for x

OpenStudy (calculusxy):

Okay. 1500 x 0.16 = 240 20x + 240 = 365 20x=125 125/20 = 6.25

OpenStudy (calculusxy):

So the answer is 6.25 right?

OpenStudy (phi):

yes. algebra is very powerful way to figure it out.

OpenStudy (calculusxy):

THANK YOU!!

OpenStudy (calculusxy):

Can you help me with some other questions? Please.

OpenStudy (phi):

a few

OpenStudy (calculusxy):

Thanks :)